Suppose the required reserve ratio is 20%. If the Fed creates $100 in bank reserves,
and the banks loan out the maximum each time and the first borrower takes the loan in
the form of cash which they walk out the door with, how much will the money supply
increase?

What if everything is the same as question 4, except instead of the Fed creating $100
in bank reserves, instead grandma takes $100 cash buried under the tree and puts it in
a checking account. How much will the money supply increase?
Answer #1

a) reserve ratio = 20%

Multiplier = 1/0.2 = 5

As the Fed created $100 in bank reserves, It means there is excess reserve of $100 in the bank and as the bank loan out the maximum each time.

He will give the $100 to the first borrower and as the first borrower takes the loan in the form of cash which they walk out the door with, so money supply will increase by only $100

b) instead grandma takes $100 cash buried under the tree and puts it in a checking account. It means deposits increase by $100

so the bank keeps the 20% and loan out the rest and he cycle will keep on.

money supply increase by 100*5 = $500